# Driving Scene Synthesis on Free-form Trajectories with Generative Prior

### v1

@@ 3.2 DriveX
where $x,y$ represent the pixel coordinates, and $d$ is the depth (Please refer to Appendix for details). Notably, $\psi$ is invertible, allowing us to unproject an image with its corresponding depth map back to 3D points given the camera pose. Consequently, we can obtain a pseudo image $\hat{I}_{\text{ren}}$ under the pose $P_{\text{ren}}$ by sampling the color in $I_{\text{rec}}$ using the warped pixel coordinates $(\mathbf{x},\mathbf{y},\mathbf{d})$ from unwarped $I_{\text{ren}}$ to $I_{\text{rec}}$, as expressed by Eq. 8.
$(\mathbf{x},\mathbf{y},\mathbf{d})=\psi(\psi^{-1}(I_{\text{ren}},D_{\text{ren}}|P_{\text{ren}})|P_{\text{rec}}),$ (8)
where $D_{\text{ren}}$ is the rendered depth corresponding to $\hat{I}_{\text{ren}}$. Since it is hard to guarantee the pixel-wise correspondence between $I_{\text{ren}}$ and $\hat{I}_{\text{ren}}$, we choose SSIM to evaluate similarity at the patch-structure level. The unreliability mask $\mathcal{M}$ is then obtained by applying a threshold $\tau$ to SSIM score:
@@
Since the warped image provides a rough overview about the novel trajectory, making it unnecessary to start denoising from scratch when we refine the videos using Eq. 5. Hence, we perturb $\hat{I}_{t}$ by adding Gaussian noise at some noise level. This consideration not only reduces unnecessary time costs but also avoids color shifts by preserving accurate low-frequency components in original video.
EmerNeRF [36] PVG [4] StreetGaussian [35] DriveX (Ours)
Recorded image DriveDreamer4D [40] DriveX (Ours)
@@
Figure 3: Qualitative comparisons of novel trajectories on Waymo. Top: For each method, images are rendered to the left of the recorded trajectory by $3m$. Bottom: Our results are consistent with the recorded image compared to DriveDreamer4D [40].
Iterative refinement
We have detailed how to use generative prior to solve a video inverse problem to improve the reconstruction quality. In turn, the generation can also benefit from these improvements, as it is conditioned on $\mathcal{G}_{t}$ whose enhancements could provide more accurate information for generation. This reciprocal enhancement motivates an iterative refinement. As illustrated in Algorithm 1, we first warm up the optimization for $T_{0}$ steps using conventional reconstruction method [35] on recorded trajectory, and then store the refined video $\mathcal{V}_{t,\text{refine}}^{\prime}$ as a buffer, which will be updated every $K$ steps. Due to the additional time demands of running video diffusion model, it is inefficient to set $K=1$, i.e., the buffer is updated every step. Please refer to Sec. 4.3 for the trade-off between different $K$. Since the generated contents are conditioned on Gaussian model $\mathcal{G}_{t}$, they are constrained with the recorded scene, thus ensuring the consistency between reconstruction and generation.
Novel trajectory sampling
Another important design space is how to sample the novel trajectory $\mathcal{T}^{\prime}=\left\{P_{i}^{\prime}\right\}_{i=0}^{F}$. The ideal trajectory should balance the following two requirements: first, maximizing the utility of the generative model, i.e., the generated novel views should contribute as much as possible to enhance reconstruction quality; second, minimizing inconsistency between the recorded and generated images. Therefore, we adopt a panning camera trajectory that starts from a recorded front view $P^{\prime}_{0}=[R_{0}|T_{0}]$ and gradually shifts laterally:
$P^{\prime}_{i}=[R_{0}|\frac{i}{F}s\mathbf{v}+T_{0}],$ (10)
where $\mathbf{v}\in\mathbb{R}^{3}$ is the shifting direction, and $s\in\mathbb{R}$ controls the maximum shifting length. This design not only provides novel views beyond the recorded trajectory, but also allows video generation model to extract detailed references from the initial frame. At the early stage of optimization, rendering quality may deteriorate as the camera diverges from the recorded trajectory, which offers limited guidance for video generation. To address this issue, we initially use a small shifting length $s$ and progressively extend it during optimization, ensuring our method’s stability and robustness across diverse scenes.
±$\pm 1m$$\pm 2m$$\pm 3m$ Inference speed
@@
EmerNeRF [36] 0.1676 0.5149 91.40 0.1372 0.4671 111.67 0.1369 0.4198 132.48 0.12
$S^{3}$aussian [12] 0.1423 0.5301 70.71 0.0865 0.4781 100.29 0.0430 0.4341 126.22 8.4
PVG [4] 0.1335 0.5406 71.63 0.0632 0.4686 102.92 0.0365 0.4267 137.12 48
StreetGaussian [35] 0.2002 0.5945 49.44 0.1668 0.5793 80.05 0.1385 0.5342 105.37 34
DriveX (Ours) 0.2065 0.5949 48.65 0.1892 0.5827 76.44 0.1755 0.5427 95.60 34
Table 1: Novel trajectory evaluation on Waymo. We compare images rendered in novel trajectories with different shift lengths, where the IoU is computed between the ground truth and those detected by TwinLiteNet [3]. Once trained, our model solely relies on the Gaussians for inference. The inference speeds of all methods are measured on the NVIDIA RTX A6000 GPU.
Gaussian model Lane change Acceleration Deceleration Average
[…]
DriveDreamer4D [40] PVG [4] 0.428 53.00 0.411 53.10 0.421 53.78 0.420 53.29
DriveX (Ours) PVG [4] 0.474 58.53 0.563 65.36 0.611 66.97 0.549 63.62
DriveX (Ours) StreetGauss. [35] 0.516 62.77 0.569 65.14 0.599 66.23 0.561 64.71
Table 2: Comparision with DriveDreamer4D [40] under its evaluation setting. For fair comparison, we report the same metric proposed in [40] measured on the same sequences, and same Gaussian model [4] with the best reported results in [40].
3.3 Application on generated videos
In addition to real-world scenes, the driving videos used for reconstruction can also be created by a video generator [7], which explores new possibilities for the automatic creation of diverse virtual driving environments. However, the AI-generated videos present more severe challenges because of the lack of corresponding LiDAR depth and accurate pixel-to-pixel matching between different frames. These challenges further necessitate the novel view supervision provided by the generative priors.
Therefore, we devise an additional recipe for the robust optimization of the AI-generated videos. Specifically, since the generated videos cannot be precisely aligned with the conditioned camera parameters, we employ a feed-forward approach [39] to estimate the camera parameters and corresponding depths for each frame. For scenes containing moving vehicles, we manually annotate their tracklets as the initialization for the local coordinate frames. During optimization, we fine-tune the camera pose to mitigate potential prediction errors. In the computation of unreliability, as the generated video is monocular, we use the nearest previous frames that can include the current shifted view content as the auxiliary source images in Eq. 8 to ensure well-defined unreliability scores for all areas of the novel view.
@@ 4.1 Experimental setup
For evaluation of the improvement for real-world reconstruction, we conduct experiments on 12 selected sequences comprising surrounding videos and synchronized LiDAR point clouds from Waymo Open Dataset [26]. Official tracklets are used to crop the point cloud of dynamic traffic participants and initialize their per-frame pose. For fair comparison, all images are downsampled into $640\times 920$ in both training and testing. To test the applicability of DriveX in generated driving videos, we assess our method on 12 driving scene videos produced by video diffusion models [7]. Each video is generated given the initial frame from nuScenes validation set as the condition.
Metrics
@@
To reasonably benchmark the improvement on the existing reconstruction-based street view synthesis pipeline, we quantitatively compared DriveX with several representative baselines on views from the novel trajectory. Since the ground truth sensor data on these trajectories is unavailable, it is infeasible to directly assess the pixel-level fidelity. Therefore, we devised a novel benchmark to comprehensively evaluate the synthesis results at these views. Specifically, the Fréchet Inception Distance (FID) [9] between synthesized views in novel trajectories and the captured images from the original trajectories is employed to assess the realism of synthesized images at the distribution level. Moreover, we report Lane IoU and vehicle AP to evaluate the distinguishability and fidelity of two safety-critical high-level traffic components, i.e., road lanes and vehicles, in novel trajectory viewpoints. Refer to the appendix for detailed calculations of these metrics.
Implementation details
@@
For reconstruction of real video where calibrated data and LiDAR are available, we first warm up the optimization with Gaussian model [35] for 50,000 steps and then integrate generative priors for following 30,000 steps. For novel trajectory videos, the initial view $P_{0}^{\prime}$ of a novel trajectory $\mathcal{T}^{\prime}$ is selected from every 3 recorded forward cameras, with shifting direction $\mathbf{v}$ oriented to the left or right, and shifting length $s$ ranging from $2m$ to $4m$. During the iterative refinement, the buffer for storing refined videos is updated every $K=2000$ steps. We use ViewCrafter [37] as our generative prior. The refine strength (i.e. the level of noise adding to images) is set to 0.6. The threshold $\tau$ in Eq. 9 is set to 0.65. To alleviate the conflicts between recorded data and generated results on moving cars, their supervision strength is reduced to 0.2. For generated videos, we first estimate the camera pose [39] and predict the depth and then perform similar process.
4.2 Comparison with state of the art
Results on real-world videos
@@ 4.2 Comparison with state of the art
For comparisons on real-world reconstruction, we primarily focused on improvements in substantial novel views beyond the recorded trajectory. To this end, we synthesized novel driving videos along trajectories shifting from the recorded trajectories by ±$\pm 1m$, ±$\pm 2m$ and ±$\pm 3m$. Table 1 demonstrate our significant improvement on FID, AP and IoU, surpassing the leading StreetGaussian [35] with a 9.3% decrease in FID, 1.6% increase in AP and 21.1% increase in IoU under the largest $3m$ shifting length. These enhancements are visually corroborated in Fig. 3, where safety-critical scene elements can be well synthesized in novel views by our methods. It can be observed that by integrating generative priors, our approach demonstrates robust rendering quality under large trajectory deviations, resulting in more realistic driving environments. Notably, once trained, our model solely relies on the Gaussians for inference, enabling efficient rendering speed.
For fair comparison with DriveDreamer4D [40] which is not open-source, we adopt the same experiment setting (sequences, metric and its best reported Gaussian counterpart [4]). The results are presented in Tab. 2 and bottom of Fig. 3. Our model markedly outperforms DriveDreamer4D with variety of Gaussian models on all metrics, visual quality, and the consistency of traffic components.

### v2

@@ 3.3 Inpainting-based video restoration
where $x,y$ represent the pixel coordinates, and $d$ is the depth (Please refer to Appendix B for details). Note that the refined video starts from a front view on the recorded trajectory, which is selected as the “closest” image for all refined views in the video. Notably, $\psi$ is invertible, allowing us to unproject an image with its corresponding depth map back to 3D points given the camera pose. Consequently, we can obtain a pseudo image $\hat{I}_{\text{ren}}$ under the pose $P_{\text{ren}}$ by sampling the color in $I_{\text{rec}}$ using the warped pixel coordinates $(\mathbf{x},\mathbf{y},\mathbf{d})$ from unwarped $I_{\text{ren}}$ to $I_{\text{rec}}$, as expressed by Eq. 9.
$(\mathbf{x},\mathbf{y},\mathbf{d})=\psi(\psi^{-1}(D_{\text{ren}}|P_{\text{ren}})|P_{\text{rec}}),$ (9)
where $D_{\text{ren}}$ is the depth map corresponding to $\hat{I}_{\text{ren}}$ rendered by $\alpha$-blending each Gaussian center’s depth. Since it is hard to guarantee the pixel-wise correspondence between $I_{\text{ren}}$ and $\hat{I}_{\text{ren}}$, we choose SSIM to evaluate similarity at the patch-structure level. The unreliability mask $\mathcal{M}$ is then obtained by applying a threshold $\tau$ to SSIM score:
@@
Since the warped image provides a rough overview of the novel trajectory, making it unnecessary to start denoising from scratch when we refine the videos using Eq. 7. Hence, we perturb warped videos by adding Gaussian noise at some noise level. This not only reduces unnecessary time costs but also avoids color shifts by preserving accurate low-frequency components in original video. Because autonomous vehicles are typically equipped with surround-view cameras, we include side view on recorded trajectory in the same frame as the auxiliary source images, which is also warped using equation (9) to the regions that are invisible by the front view to ensure well-defined unreliability masks for all areas of the novel view.
Since the sparse depth ground truth in the novel view can be derived from LiDAR measurements, the Gaussian model performs better in depth rendering than in color rendering under novel view, even in the presence of overfitting. This allows the warped images to serve as a reliable anchor for indicating the reliability.
Projected LiDAR as auxiliary condition Additionally, we also utilize colored LiDAR projections as an oracle to guide the generative process. For each novel view requiring restoration, LiDAR points projected onto it are accumulated from its adjacent ±$\pm 2$ frames, where points belonging to moving objects are aligned to the rendered frame based on their tracked poses. This empirically stabilizes the optimization, especially for those challenging scenes where rendered novel views struggle to provide reliable information in the early optimization. In light of the importance of novel view depth supervision within this framework, we additionally use the accumulated LiDAR point cloud as the depth ground truth for novel views.
@@
Figure 3: Qualitative comparisons on novel trajectories. For each method, images are rendered at the novel views with lateral shift of $3m$ to the recorded trajectories.
3.4 Training of generative model
One remaining question is how to obtain a generative prior capable of handling the restoration task in Eq. 7. Due to the limited trajectory in existing driving datasets, precisely simulating degradation in novel trajectory views caused by overfitting has an inherent challenge. Fortunately, our inpainting approach can mitigate this issue by disentangling the degraded pattern from the condition, avoiding the necessity to simulate degraded rendering during training and freeing the model from identifying potential degradations.
Specifically, we employ an edge-aware strategy to generate unreliability masks for the training data. This leads to a more general task which can cover scenarios typically encountered during off-trajectory rendering. Given a RGB image, we first convert it into grayscale. Then its directional derivatives computed using Sobel filters, as the edge intensity at each pixel. The resulting intensity map is normalized to yield a probability distribution over the pixels after adding a baseline offset for background regions. Then a predefined number of pixels is randomly sampled based on the distribution. For each sampled pixel, a $3\times 3$ patch centered at the pixel is masked out to imitate the structured occlusion. Both front and side-view camera videos are used in training. Because the trajectory of the side-view camera along the moving forward trajectory is similar to the lateral shift of the front cameras adopted in our novel trajectory sampling, as suggested by [28, 4].
While there remains a gap between training and inference, it does not substantially hinder the usage. This is because the training process primarily aims at unleashing the generative capability of the diffusion model to adapt on such a masked image inpainting task, rather than fitting specific mask patterns. Actually, even [39] trained only on rendered point clouds can also serve as a competent generative prior for the crafted inpainting-based restoration task. While fine-tuning the model on driving videos can further adapt it to the style of driving scenarios, thereby significantly improving its robustness as demonstrated in Sec. 4.4.
±$\pm 0m$ (recorded) ±$\pm 1m$$\pm 2m$$\pm 3m$ FPS
@@ 3.4 Training of generative model
EmerNeRF [38] 29.60 0.8459 0.1147 0.5101 89.43 0.0881 0.4644 102.98 0.0735 0.4072 122.83 0.12
$S^{3}$aussian [11] 29.68 0.8670 0.0718 0.4499 115.67 0.0401 0.4127 131.26 0.0177 0.3812 162.03 8.4
PVG [3] 29.98 0.8627 0.1082 0.5444 51.86 0.0291 0.5102 86.93 0.0170 0.4699 123.53 48
StreetGaussian [37] 29.76 0.8738 0.2671 0.5919 51.87 0.2122 0.5848 70.66 0.1720 0.5697 93.04 34
DriveX (ours) 29.74 0.8700 0.2880 0.5932 46.91 0.2710 0.5901 65.52 0.2699 0.5712 79.78 34
Table 1: Quantitative evaluation on Waymo dataset [25]. We compare images rendered in novel trajectories with different shift lengths, where the IoU and AP are computed between the projected ground truth and those detected on synthesized views by perception models [21, 2]. The inference speeds of all methods are measured on the NVIDIA RTX A6000 GPU.
Figure 4: Qualitative comparisons with methods incorperating generative model. The results of DriveDreamer4D [44] and ReconDreamer [20] are downloaded from their project pages. The result of FreeVS [28] is obtained by running its official released code. Compared to other alternatives, our method can more effectively harness the generative prior, significantly reducing hallucinations.
@@ 4.1 Experimental setup
Datasets   We evaluate our framework and train the diffusion model on driving sequences from large-scale Waymo Open Dataset (WOD) [25]. Each sequence comprises surrounding videos and synchronized LiDAR point clouds capture at 10Hz. For training, we use all sequences from its training and validation set. The quantitative evaluation is conducted on 18 selected sequences, each containing 40 frames. Official bounding boxes are used to crop the point cloud of dynamic traffic participants and initialize their per-frame pose. For fair comparison, all images are downsampled into $640\times 960$ in both training and evaluation. More details can be found in Appendix B.
@@
Metrics   We evaluate the view synthesis quality on both recorded and novel trajectories. For recorded trajectory, we follow the common practice of reporting the PSNR and SSIM on these views. The test views are held out from every 5 frames. For novel trajectory views, since the ground truth is unavailable, it is infeasible to directly assess the pixel-level fidelity. Therefore, we devised a novel benchmark to comprehensively evaluate the synthesis results at these views. Specifically, the Fréchet Inception Distance (FID) [8] between synthesized views in novel trajectories and the captured images from the original trajectories is employed to assess the realism of synthesized images at the distribution level. Moreover, we report Lane IoU and vehicle AP (Refer to Appendix B for these calculations.) to evaluate the distinguishability and fidelity of two safety-critical high-level traffic components, i.e., road lanes and vehicles, in novel trajectory viewpoints.
@@
Implementation details   For the reconstruction with calibrated camera images and LiDAR point clouds, we first initialize the scene representation [37] by optimizing on the recorded data for $T_{0}=50,000$ steps and then integrate generative priors for following $30,000$ steps. For the generation of novel trajectory videos, the initial view $P_{0}^{\prime}$ of a novel trajectory $\mathcal{T}^{\prime}$ is selected from every 3 frames, with shifting length $s$ ranging from $2m$ to $6m$. The generative model is initialized from  [39], and then fine-tuned on WOD [25] for $15,000$ iterations with batch size of 24 using 8 RTX A6000 GPUs. During the iterative refinement, the buffer for storing refined videos is updated every $K=3000$ steps. The refine strength (i.e. the level of noise adding to images) is set to 0.6. The threshold $\tau$ in Eq. 10 is set to 0.65.
Figure 5: Ablations on (a) unreliability mask and (b) iterative refinement. (a) shows the initial restored results without (middle left) and with (middle right) unreliability mask (top right). The final rendered results exhibit significant artifacts without mask (bottom left). (b) shows the rendered (top) and restored results at iteration $T_{0}$ (middle left) and $T$ (middle right). It can be seen that single-time refinement can not completely eliminate artifacts (as shown in orange boxes), resulting in inferior final results compared to iterative refinement (bottom). Best viewed with zoom-in.
4.2 Main results
@@
In Tab. 1, we quantitatively compared DriveX with several representative baselines in terms of view synthesis quality and rendering speed. For extrapolated novel views, we synthesized driving videos laterally shifting from the recorded trajectories by ±$\pm 1m$, ±$\pm 2m$ and ±$\pm 3m$. The result demonstrates our significant improvement on FID, AP and IoU, surpassing the leading reconstruction-based counterpart [37] with a 14.3% decrease in FID and 56.9% increase in IoU under the $3m$ shifting length. These enhancements are visually corroborated in Fig. 3, where our approach demonstrates robust rendering quality under large trajectory deviations, where safety-critical scene elements can be more clearly recognized in novel views by incorporating generative priors. Notably, once trained, our model solely relies on the 3DGS for inference, enabling efficient rendering speed. For views on recorded trajectories, our method also achieves comparable performance with the baseline. The slightly lower metrics often result in perceptually negligible impact on synthesized quality. This further indicates our method can seamlessly integrate generative prior by effectively minimizing inconsistencies of generated content.
Lane change Lane shift ±$\pm 3m$
[…]
DriveDreamer4D [44] 0.495 53.42 0.340 51.32
ReconDreamer [20] 0.554 56.63 0.539 54.58
DriveX (ours) 0.620 58.60 0.567 58.29
Lane shift ±$\pm 1m$ Lane shift ±$\pm 3m$
[…]
FreeVS [28] 0.2782 0.5777 0.2723 0.5688
DriveX (ours) 0.2880 0.5932 0.2699 0.5712
Table 2: Comparison with methods using generative model. All results are averaged from 8 sequences used in  [44]. For lane change, the NTA-IoU and NTL-IoU are evaluated under setting in [44, 20], and we report the best results of [44, 20] from their original paper. lane shift is computed in the same way as our main paper. For FreeVS [28], we run its official released model and measure the results on the same sequences, note that its cropped image size does not essentially impact the results.
4.3 Comparison with methods incorporating generative prior
In previous section, we compared our methods with popular reconstruction-based approaches as the main results because it is directly built upon these reconstruction-based solutions and primarily aims to address the view extrapolation challenge for them. Although there are some concurrent works [44, 20, 28, 4, 41] sharing a similar scope with ours, most of them have not released code until our paper submission. Despite that, we try our best to compare with those allowing for a reasonable comparison and report the results in Tab. 2. The results indicate that our method performs better than other alternatives, which is also confirmed by qualitative results in Fig. 4. Compared to DriveDreamer4D [44], our approach substantially eliminates hallucinations and inconsistencies caused by insufficient constraints. Unlike generative-only approach FreeVS [28], our method maintains continuous 3D scene representation, enabling synthesized views to preserve superior detail fidelity and better consistency across frames than those only relying on sparse LiDAR signal. Moreover, our method has significantly faster synthesis speed (34 FPS vs. 0.43 FPS).
[…]
mask all 0.2153 0.6334 102.69
w/o mask 0.2193 0.6341 76.26
w/ mask 0.2263 0.6389 74.34
Table 3: Effects of unreliability mask. “mask all”, “w/o mask”, and “w/ mask” denotes the threshold $\tau$ in Eq. 10 set to 1.0, -1.0, and 0.65, respectively.
[…]
500 0.2271 0.6392 75.33 4.8×4.8$\times$4.8 ×
3,000 0.2263 0.6389 74.34 1.6×1.6$\times$1.6 ×
6,000 0.2228 0.6388 75.98 1.3×1.3$\times$1.3 ×
Inf 0.2194 0.6173 89.06 $\mathbf{1.0\times}$
Table 4: Ablations on buffer intervals.Inf” denotes the novel trajectory supervision only generated one time at iteration $T_{0}$.
4.4 Ablation study
Unreliability mask   The unreliability mask is designed to guide the generation of video diffusion by directly indicating where the model needs to be retained or repaired. From Tab. 3 and Fig. 5(a) we can observe: (i) When all regions are masked, the video diffusion model can only rely on sparse LiDAR projections as condition, leading to inferior results due to the inability to fully leverage information from the in-training Gaussian representation; (ii) When the mask is disabled, the generative model can still refine the rendered results toward the learned video distribution, but usually with too conservative refinement (it also indicates that the model does not strictly follow the given mask during inpainting); (iii) Equipped with the proposed unreliability mask, our full model achieves superior performance in both quantitative and qualitative ablation.
Iterative refinement   The interval of iterative refinement offers the trade-off between quality and efficiency. When the novel trajectory supervision only generated once time at iteration $T_{0}$ (see Fig. 5(b)), the quality of the novel view is constrained to the in-training Gaussian model, which is overfitted on the recorded trajectory. To the other extreme, we are unable to afford the time cost of running the diffusion at every step. From Tab. 4, it can be seen that even a small interval 500 leads to 4.8×4.8$\times$4.8 × training time. Hence we finally set the interval to 3,000 to strike the balance.
Figure 6: Domain-specific fine-tuning enhances the style adaptation on driving videos. Here the appearance of the black car refined by model without fine-tuning seems to lack details and has unrealistic material.
Projected LiDAR points as auxiliary oracle   Although using the masked rendered image as the only condition is sufficient to achieve satisfactory performance in most scenes, incorporating projected LiDAR points as the additional oracle can greatly improve robustness in some difficult cases. In these scenes, initial poor novel view renderings would result in large masked regions, LiDAR projection can provide direct guidance for precisely restoring them. This resulted enhancement is reflected by the quantitative comparison in Tab. 5.
[…]
render LiDAR ft.
0.2092 0.6257 90.30
0.2123 0.6334 86.30
0.2153 0.6225 75.18
0.2263 0.6389 74.34
Table 5: Ablations on the components of generative model. “w/ render” means using masked novel view image as condition. “w/ LiDAR” denotes adopting LiDAR projection as condition. “w/ ft.” represents training diffusion model on driving video data [25].
Fine-tuning generative model on driving videos   The comparison between the $2^{nd}$ and $4^{th}$ row of Tab. 5 demonstrates that the diffusion model fine-tuned on in-domain driving videos outperforms that trained solely on general-purpose 3D datasets [39], especially in FID. This improvement comes from the generative model’s adaptation to the style of real driving scenes. As shown in Fig. 6, the counterpart without domain-specific fine-tuning tends to generate cartooned foreground vehicles without detailed texture. Therefore, beyond the improvement in quantitative metrics, it also simplifies our pipeline by removing the need for some hand-designed processing, such as lowering the refine weight for foreground vehicles.
5 Conclusion
In this paper, we present a novel framework that integrates generative priors into the reconstruction of the driving scene from single-trajectory recorded videos, addressing limitations in extrapolating novel views beyond a recorded trajectory. To reasonably leverage the video generative model to provide consistent novel view supervision, we construct a tailored inpainting-based video restoration task to refine the novel view renderings for optimizing the street scene representation. The experiments demonstrate that the proposed method significantly improves the synthesis quality for novel trajectories. These results pave the new way for driving scene synthesis on free-form trajectories.
Acknowledgments
This work was supported in part by National Natural Science Foundation of China (Grant No. 62376060).
Appendix
Appendix A Discussion on limitations
Beyond demonstrating the efficacy of the proposed method in synthesizing driving views along arbitrary trajectories, this paper also acknowledges several limitations. First, incorporating a diffusion-based generative model significantly increases reconstruction time. But such complexity might be unnecessary, as the required generative diversity is substantially constrained under our inpainting formulation. suggesting that a more lightweight model can be employed in the future. Second, our method currently employs a simple panning camera scheme to sample novel trajectories to be restored. However, this fix design may not be optimal for all scenarios, and leave some hand-designed components in the pipeline. Future work could consider to adaptively explore the optimal novel trajectories in optimization.
Appendix B More implementation details
Metric details
Previous driving scene reconstruction methods [40, 7, 12] typically utilize PSNR, SSIM [31] and LPIPS [43] for evaluate image level similarity with ground truth. However, these metrics are inapplicable for evaluating free-form trajectories due to the absence of corresponding ground-truth images for novel trajectories.
Therefore, inspired by [28, 44, 20], we propose a novel set of metrics to evaluate the recognizability of high-level traffic elements in synthesized views, based on ground truth annotations in 3D space. For vehicles, we projected their ground truth 3D bounding boxes onto the novel views to obtain corresponding 2D bounding boxes. Since IoU lacks the precision granularity needed to understand performance across various intersection over union threshold. Therefore, we calculate the Average Precision (AP[50:95]) between the projected ground truth and those detected by Faster R-CNN [21], which provides a more nuanced assessment than IoU.
For road lanes, NTL-IoU adopted in  [44, 20] calculates the mIoU between the lanes on synthesized views detected by [2] and the ground truth, averaging both foreground lanes and backgrounds. However, including the background can inflate accuracy scores due to the easy detection of non-lane areas. Therefore, following standard lane segmentation evaluation [2], we focus specifically on IoU with only the foreground to better evaluate lane synthesis quality.
Warping operation $\phi$ in Eq. (8) of main paper
$(x_{*},y_{*},d)^{\top}$ $=Kp_{\text{camera}}$ (12)
$(x,y)$ $=(\frac{x_{*}}{d},\frac{y_{*}}{d})$ (13)
Validation sequences
There are 18 curated driving sequences from WOD [25] used for qualitative and quantitative evaluations in our paper. The names of all segments are listed below:
  • 10359308928573410754_720_000_740_000
  • 11450298750351730790_1431_750_1451_750
  • 12496433400137459534_120_000_140_000
  • 15021599536622641101_556_150_576_150
  • 16767575238225610271_5185_000_5205_000
  • 17860546506509760757_6040_000_6060_000
  • 3015436519694987712_1300_000_1320_000
  • 6637600600814023975_2235_000_2255_000
  • 10444454289801298640_4360_000_4380_000
  • 10588771936253546636_2300_000_2320_000
  • 10625026498155904401_200_000_220_000
  • 11017034898130016754_697_830_717_830
  • 1191788760630624072_3880_000_3900_000
  • 11928449532664718059_1200_000_1220_000
  • 14810689888487451189_720_000_740_000
  • 4414235478445376689_2020_000_2040_000
  • 6242822583398487496_73_000_93_000
  • 7670103006580549715_360_000_380_000
Unless stated otherwise, all quantitative ablation experiments are conducted on segments 15021599536622641101_556_150_576_150 and 3015436519694987712_1300_000_1320_000, while the results are evaluated on novel trajectories with lateral shift $3m$.
Appendix C Additional ablations
[…]
0.4 0.2184 0.6202 96.21 $\mathbf{1.0\times}$
0.6 0.2263 0.6389 74.34 $\times1$
0.8 0.2192 0.6259 81.55 1.2×1.2$\times$1.2 ×
1.0 0.2113 0.6293 106.01 1.3×1.3$\times$1.3 ×
Table 6: Effect of the noise level. The noise level $l$ denotes adding the noise corresponding to the timestep $t=\lfloor lN\rfloor$, where $N$ is the total step needed for denoising from scratch.
[…]
PVG [3] 0.0200 0.5401 118.41
PVG [3] 0.1872 0.6197 91.85
StreetGaussian [37] 0.1217 0.6124 103.08
StreetGaussian [37] 0.2263 0.6389 74.34
Table 7: Ablation study on different Gaussian model.
[…]
w/o progress. 0.2204 0.6372 75.41
w/o warp 0.2196 0.6357 105.77
Full model 0.2263 0.6389 74.34
Table 8: Ablations on other design choices. “w/o progress” denotes that the shift length is fixed at $4m$ throughout the optimization. “w/o warp” means that the denoising is started from the noisy rendered images rather than warped views.
Refine strength
When we refine the rendered videos using video diffusion, we also study the effect the refine strength (level of noise added to the images). In Tab. 6, we show the metrics for novel trajectory synthesis (IoU, AP, and FID), along with the total optimization time. With a low refine strength, the refined image adheres to the artifact-heavy rendered image, resulting in minimal improvement. On the other hand, a high refine strength yields better novel views but may not keep fidelity to the original scene. Additionally, higher refine strength requires more denoising steps, increasing the training time. Consequently, refine strength of 0.6 is found to be the best balance.
Different Gaussian model
Although we adopt static 3D Gaussian splatting [37] as the default scene representation in previous experiments, our method is also capable of synergy with different choices. To demonstrate the versatility of DriveX, we also evaluate our method on a representative dynamic Gaussian representation for driving view synthesis [3]. From Tab. 7, we can observe that our strategy consistently reduces the FID by a large margin by substantially reducing artifacts in novel views, and effectively recovering street lanes and cars (i.e. IoU, AP) that are entirely indiscernible in the baseline.